# Recovery speed

### v1

@@ 7 Experiments
In this experiment, we aim to verify two aspects of our proposed method. Specifically, we will verify the following:
(A) $k$-dependency of $\bar{T}_{R}(k)$: Verify that increasing $k$ in $\bar{T}_{R}(k)$ leads to more accurate recovery time estimates.
(B) Control $T_{R}$: Verify that adjusting $\varepsilon$ enables the tuning of $T_{R}$, demonstrating a trade-off between recovery time and inference loss.
Detailed experimental settings are given in Appendix C. In Appendix D, we compare our resilience training method and adversarial training.
@@ 7.1.1 Experimental Setup
We employed a simplified model using an LSTM network with randomly initialized weight parameters. We applied step input to this model, introduced instantaneous noise, and observed the time required for inference recovery (for a visual representation of the recovery time measurement, see Figure 1).
@@ 7.1.2 Results and Analysis
Figure 2: k-dependency of $\bar{T}_{R}(k)$. Each point represents a pair of the correlation coefficient and estimation error between $\bar{T}_{R}(k)$ and $T_{R}$ for $k=0,1,\ldots,20$ (averaged over $20$ models). Edges connect points corresponding to consecutive $k$ values. As $k$ increases, the points move from the upper left to the lower right of the figure.
Figure 3: Accuracy of recovery time estimation. Gray dots represent $T_{R}$ for individual models. Orange and blue dots indicate $\bar{T}_{R}(k)$ for each model. The red line represents the $\delta$ISS sufficient condition boundary. According to Theorem 4.2, models to its left are guaranteed $\delta$ISS.
Figure 4: Recovery time of simplified model. The blue and orange lines depict the LSTM outputs for the normal input without noise and the anomalous input subjected to instantaneous noise, respectively. Vertical lines denote the recovery times.
We calculated $T_{R}$ and $\bar{T}_{R}(k)$ $(k=0,1,\cdots,20)$ for randomly initialized $20$ models, deriving estimated average test errors and correlation coefficients. As shown in Figure 2, the estimation test error consistently decreases as $k$ increases, eventually converging at around $k=20$ steps. Furthermore, the correlation coefficient exhibits a generally increasing trend as $k$ increases, excluding fluctuations for large $k$, which indicates a strong relationship between the true and our estimated values.
Figure 3 compares the estimation accuracy for $k=0$ and $k=20$, demonstrating improved accuracy with a larger $k$. The red line indicates the $\delta$ISS guarantee boundary from Theorem 4.2. Increasing $k$ from 00 to $20$ expands the range of models guaranteed to be $\delta$ISS, as shown by more data points to the left of the red line for $k=20$.
To illustrate this improvement in a specific case, we focus on one of these models. As evident in Figure 4, the $\bar{T}_{R}(k)$ value for $k=20$ is smaller than that for $k=0$. Furthermore, this value is closer to the true $T_{R}$. As a result, using a larger $k$ value enables a more accurate assessment.
@@
7.2 Experiment $\rm{I}\rm{I}$: Two-Tank System
In the previous subsection, we verified conditions (A) using a simplified model. Here, we demonstrate the effectiveness of our approach in a more practical scenario using the two-tank system, a recognized benchmark in control experiments Jung et al. (2023); Schoukens & Noël (2017). We focus on demonstrating condition (B), which could not be experimentally verified using simplified models.
@@ 7.2.1 Experimental Setup
System: The two-tank system is described by the following differential equations Schoukens & Noël (2017):
@@
where $u$ is the control input, $h_{1}$ and $h_{2}$ are the tank levels, $g$ is the gravitational acceleration and $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ are hyperparameters depending on the system properties. The LSTM model is designed to predict the water level at the next time step based on the current water level and control input. The current water level is obtained from sensors, which are assumed to be subject to sensor noise. Formally, we define the input-output relationship of the LSTM as $x(t)=(u(t),h(t)+w(t))$ and $y(t)=h(t+1)$ where $w\sim\mathcal{N}(0,{0.01}^{2})$.
Baseline: To evaluate our proposed recovery time estimation method and adjustment technique, we prepared a baseline model: a standard model trained without considering resilience. As shown in Figure 5, the baseline model closely approximates the target system for inference.
Figure 5: Baseline of two-tank system. Top: The blue line depicts the resulting tank water level $h_{1}$ while the orange line illustrates the LSTM prediction. Bottom: control input signal.
@@ 7.2.2 Results and Analysis
First, regarding (A), our experiments corroborated the wide-sense monotonic decrease in the $k$-dependency of $\bar{T}_{R}(k)$, as stated in Theorem 5.3 and Remark  5.4, and as illustrated in Figure 6. Although not contradicting the theorems, the magnitude of this decay was notably less pronounced compared to the results obtained from the simplified model simulations in Section 5.1.
Next, concerning (B), we observed a trade-off between recovery time and MAE loss when comparing the baseline model and our models, as illustrated in Figure 6. Additionally, there is an expected trade-off relationship between $T_{R}$ and MAE loss as $\varepsilon$ varies. Specifically, as $\varepsilon$ increases from $0.0$ to $0.5$, we observe a trend of increasing MAE loss and decreasing recovery time. This trend aligns with our theoretical predictions in Section 6.2.
@@
Figure 6: Trade-off controlled by $\varepsilon$. The red dot represents the baseline value. For each $\varepsilon$ value, gray, orange, and blue dots represent $(\text{MAE},T_{R})$, $(\text{MAE},\bar{T}_{R}(0))$ and $(\text{MAE},\bar{T}_{R}(20))$ pairs respectively, where MAE is the loss calculated on test data. Edges connect results for adjacent $\varepsilon$ values, illustrating the progression of outcomes as $\varepsilon$ changes.
@@ 8 Discussion
  • Theoretical and data-independent guarantees may not always reflect actual data distributions. Therefore, we strongly recommend complementing our approach with data-driven quality assurance methods for practical implementation. It is also important to note that applying our method to an unstable system can be counterproductive.
@@
  • Although our current research focused on the resilience of LSTM modules in control systems, comprehensive quality assurance for control systems remains a challenging issue. Refining studies that have demonstrated the relationship between LSTM stability and control system stability Terzi et al. (2021) could be a key approach. By more precisely analyzing the calculations used to prove control system stability in these studies, it may be possible to quantitatively determine the convergence time of the state of the system to the target value.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Other Related Work
Finite time stability (FTS) Bhat & Bernstein (2000) rigorously discusses convergence time, characterized by a settling-time function. Finite-time input-to-state stability (FTISS), combining FTS and ISS, enables robust control against disturbances He et al. (2022). Recent studies Liang & Liang (2023); Mo et al. (2023) propose Lyapunov functions satisfying FTISS, eliminating the need to directly consider the settling-time function.
However, deriving sufficient FTS conditions involves strict model constraints, established only for specific models like Hopfield networks Michalak & Nowakowski (2017) and Clifford-valued RNNs Shen & Meng (2023), but not for conventional RNNs or LSTMs. While achieving FTS for LSTMs would theoretically allow for precise recovery time evaluation, we argue that from a practical standpoint, considering numerical errors and rounding, it suffices for output differences to fall below a threshold rather than strictly reaching equilibrium. Thus, we opted for $\delta$ISS instead of FTS.
@@ Appendix B Proofs
From the definition in Section 4.1, $G^{(l)}_{*}(\eta^{(l)}(k-1)),G^{(l)}_{c}(\eta^{(l)}(k-1))\geq 0$. Since $\sigma$ and $\phi$ are strictly monotonically increasing functions, we have $11/2=\sigma(0)\leq\overline{\sigma}^{(l)}_{*}(k)<$ 1 and $10=\phi(0)\leq\overline{\phi}^{(l)}_{c}(k)<$ 1. Therefore, $\overline{\sigma}^{(l)}_{i}(k)\overline{\phi}^{(l)}_{c}(k)\geq 0$ and $0<1-\overline{\sigma}^{(l)}_{f}(k))\leq 1/2$ are obtained, then,
$\eta(k)=\phi\left(\frac{\overline{\sigma}^{(l)}_{i}(k)\bar{\phi}_{c}(k)}{1-\overline{\sigma}^{(l)}_{f}(k)}\right)\overline{\sigma}^{(l)}_{o}(k)\geq\phi\left(\frac{0}{2}\right)\frac{1}{2}=0.$ (41)
@@
Firstly, consider the case of $k=0$. From the above inequalities, and the fact that $\sigma(x)\to\infty$ as $x\to\infty$, and that $10<\overline{\sigma}^{(l)}_{o}(0)<$ 1, we have
@@
Since $\sigma$ and $\phi$ are strictly monotonically increasing functions,
$\overline{\sigma}^{(l)}_{*}(1)$ $=\sigma(G_{*}(\eta^{(l)}(0))\leq\sigma(G_{*}(\eta^{(l)}(-1))=\overline{\sigma}^{(l)}_{k}(0)$ (48)
$\overline{\phi}^{(l)}_{c}(1)$ $=\phi(G_{c}(\eta^{(l)}(0))\leq\phi(G_{c}(\eta^{(l)}(-1))=\overline{\phi}^{(l)}_{c}(0)$ (49)
Therefore, $\overline{\sigma}^{(l)}_{*}(1)\leq\overline{\sigma}^{(l)}_{*}(1)$ and $\overline{\phi}^{(l)}_{c}(1)\leq\overline{\phi}^{(l)}_{c}(0)$. Applying these inequality for $\overline{c}^{(l)}(1)$ and $\eta^{(l)}(1)$,
@@
$\quad+\sigma(W^{(l)}_{i}x^{(l)}(t)+U^{(l)}_{i}h^{(l)}(t)+b^{(l)}_{i})_{(j)}\times\phi(W^{(l)}_{c}x^{(l)}(t)+U^{(l)}_{c}h^{(l)}(t)+b^{(l)}_{c})_{(j)},$ (61)
@@ B.2 Approximaion of $\eta^{(l)}(\infty)$
Definition B.3 (ISS, Sontag 1990; Jiang & Wang 2001).
@@
Remarking the invariant set in Proposition 4.1, the sufficient condition for ISS of LSTM is relaxed (in the case of $k\geq 1$) than (Terzi et al., 2021, Theorem 1, which is the case of $k=0$) as follows.
@@
From (3) and the fact Lipschitz constant of $\phi$ is 1, we have
@@
for any $k\in\mathbb{Z}_{\geq 0}$ and $l\in[L]$. Similarly, (4) imply that
@@
for $l=1,\ldots,L-1$, where $\llcorner A\lrcorner$ denotes the lower horizontal vector $(A_{(2,1)},A_{(2,2)})$ for a matrix $A=\begin{bmatrix}A_{(1,1)}&A_{(1,2)}\\ A_{(2,1)}&A_{(2,2)}\end{bmatrix}$. Applying the relation (6) and inequality (103) to the second term of (102) repeatedly, we estimate that
@@
$\quad+\left\|{\sum_{\tau_{L}=1}^{t}A_{s}^{(L)}(k)^{t-\tau_{L}}a_{x}^{(L)}(k)\sum_{\tau_{L-1}=1}^{\tau_{L}}\llcorner A_{s}^{(L-1)}(k)^{\tau_{L}-\tau_{L-1}}\lrcorner a_{x}^{(L-1)}(k)\left\|{h^{(L-2)}(\tau_{L-1})}\right\|}\right\|,$ (106)
@@
where $M_{L}(k),M^{\prime}_{L}(k)$, and $M^{\prime\prime}_{L}(k)$ are constants independent of $t$ and $10<\rho_{L}(k),\rho^{\prime}_{L}(k)<$ 1. Therefore, if $\rho(A_{s}^{(l)}(k))\in(0,1)$ for any $l\in[L]$, the multi-layer LSTM is ISS.
We note that $\mu^{(l)}$ can be expressed by directly calculation as
@@
$(A_{s}^{(l)}(k)-\rho(A^{{(l)}}_{s}(k+1))I)v=A^{(l)}_{s}(k)v-\rho(A^{{(l)}}_{s}(k+1))v=(A^{(l)}_{s}(k)-A^{(l)}_{s}(k+1))v$ (131)
is non-negative. Since all eigenvalues of $A^{{(l)}}_{s}(k)$ less than $\rho(A^{{(l)}}_{s}(k+1))$, then $\rho(A^{{(l)}}_{s}(k+1))I-A^{(l)}_{s}(k)$ is a regular matrix, and which inverse matrix can be expanded as
@@
Therefore, $(\rho(A^{{(l)}}_{s}(k+1))I-A^{(l)}_{s}(k))^{-1}$ is positive matrix. Hence, $v$ is non-positive from (131). Because $v$ is an eigenvector, $v\neq 0$. Thus, we conclude that $\rho(A^{{(l)}}_{s}(k))\geq\rho(A^{{(l)}}_{s}(k+1))$. ∎
@@
From (3) and (4), we have
@@
Since the Lipschitz constant of $\sigma$ is $1/4$, we obtain that
@@
$\begin{pmatrix}\left\|{c_{1}^{(l)}(t+1)-c_{2}^{(l)}(t+1)}\right\|\\ \left\|{h_{1}^{(l)}(t+1)-h_{2}^{(l)}(t+1)}\right\|\end{pmatrix}$ (159)
$\leq\begin{bmatrix}\overline{\sigma}_{f}^{(l)}(k)&\alpha_{s}^{(l)}(k)\\ \overline{\sigma}_{o}^{(l)}(k)\overline{\sigma}_{f}^{(l)}&\alpha_{s}^{(l)}(k)\overline{\sigma}_{o}^{(l)}(k)+\frac{1}{4}\phi\left(\overline{c}^{(l)}(k)\right)\left\|{U_{o}^{(l)}}\right\|\end{bmatrix}\begin{pmatrix}\left\|{c_{1}^{(l)}(t)-c_{2}^{(l)}(t)}\right\|\\ \left\|{h_{1}^{(l)}(t)-h_{2}^{(l)}(t)}\right\|\end{pmatrix}$ (160)
@@
where $\alpha_{s}^{(l)}(k)=\frac{1}{4}\overline{c}^{(l)}(k)\left\|{U_{f}^{(l)}}\right\|+\overline{\sigma}_{i}^{(l)}(k)\left\|{U_{c}^{(l)}}\right\|+\frac{1}{4}\overline{\phi}_{c}^{(l)}(k)\left\|{U_{i}^{(l)}}\right\|,\alpha_{x}^{(l)}(k)=\frac{1}{4}\overline{c}^{(l)}(k)\left\|{W_{f}^{(l)}}\right\|+\overline{\sigma}_{i}^{(l)}(k)\left\|{W_{c}^{(l)}}\right\|+\frac{1}{4}\overline{\phi}_{c}^{(l)}(k)\left\|{W_{i}^{(l)}}\right\|$.
@@
$\left\|{s_{1}(t)-s_{2}(t)}\right\|$ $\leq\mu\lambda_{1}(A_{s}(k))^{t}\left\|{s_{1}(0)-s_{2}(0)}\right\|+\left\|{(I-A_{s}(k)^{t})(I-A_{s}(k))^{-1}a_{x}(k)}\right\|\left\|{x_{1}-x_{2}}\right\|_{2,\infty}.$ (163)
@@
$\left\|{s_{1}^{(L)}(t)-s_{2}^{(L)}(t)}\right\|$ (164)
$\leq\mu^{(L)}\rho(A_{s}^{(L)}(k))^{t}\left\|{s_{1}^{(L)}(0)-s_{2}^{(L)}(0)}\right\|$ (165)

### v4

@@ 7 Experiments
In this section, we verify the following two properties of our proposed method:
  • (A)
    Increasing $k$ in $\bar{T}_{R}(k)$ leads to more accurate recovery time estimates.
  • (B)
    Adjusting $\varepsilon$ enables the tuning of $T_{R}$, and there is a trade-off between recovery time and inference loss.
Detailed experimental settings are given in Appendix C.
We also compare our resilience training method and adversarial training in Appendix D.
@@ 7.1.1 Experimental Setup
We employed a simplified model using an LSTM network with randomly initialized weight parameters. We applied step input to this model, introduced instantaneous noise, and observed the time required for inference recovery (for a visual representation of the recovery time measurement, see Figure LABEL:fig:image_recovery_time).
@@ 7.1.2 Results and Analysis
We calculated $T_{R}$ and $\bar{T}_{R}(k)$ $(k=0,1,\cdots,20)$ for randomly initialized 202020 models, deriving estimated average test errors and correlation coefficients. As shown in Figure 1, the estimation test error consistently decreases as $k$ increases, eventually converging at around $k=20$ steps. Furthermore, the correlation coefficient exhibits a generally increasing (except for a temporary decrease) trend as $k$ increases, which indicates a strong relationship between the true and our estimated values.
Figure 2 compares the estimation accuracy for $k=0$ and $k=20$, demonstrating improved accuracy with a larger $k$. The left side of the red line is the area where $\delta$ISS is guaranteed by Theorem 4.2. Increasing $k$ from 0 to 202020 decreases $\rho(A(k))$ by 32% on average, then all data points are included in the $\delta$ISS guaranteed area.
To illustrate this improvement in a specific case, we focus on one of these models. As evident in Figure LABEL:fig:recovery_toymodel, the $\bar{T}_{R}(k)$ value for $k=20$ is smaller than that for $k=0$. Furthermore, this value is closer to the true $T_{R}$. This indicates that using a larger $k$ value enables a more accurate assessment.
@@
Figure 1: k-dependency of $\bar{T}_{R}(k)$. Each point represents a pair of the correlation coefficient and estimation error between $\bar{T}_{R}(k)$ and $T_{R}$ for $k=0,1,\ldots,20$ (averaged over 202020 models). Edges connect points corresponding to consecutive $k$ values. The figure shows that, as $k$ increases, the correlation coefficient tends to increase, and the estimation error monotonically decreases.
Figure 2: Accuracy of recovery time estimation. Gray dots represent $T_{R}$ for each model. Orange and blue dots indicate $\bar{T}_{R}(k)$ in $k=0$ and $k=20$ for each model. The red line represents the $\delta$ISS sufficient condition boundary. The models in the left side area of the red line are guaranteed $\delta$ISS by Theorem 4.2. Increasing $k$ from 0 to 202020 decreases $\rho(A(k))$ by 32% on average over 20 models. This figure indicates that increasing $k$ guarantees more models to be $\delta$ISS.
7.2 Experiment $\rm{I}\hskip-1.0pt\rm{I}$: Two-Tank System
Here, we demonstrate the effectiveness of our approach from the perspective of (B) in a more practical scenario using the two-tank system, a recognized benchmark in control experiments (Jung et al.,, 2023; Schoukens and Noël,, 2017).
@@ 7.2.1 Experimental Setup
System
The two-tank system is described by the following differential equations (Schoukens and Noël,, 2017):
@@
where $u$ is the control input, $h_{1}$ and $h_{2}$ are the tank levels, $g$ is the gravitational acceleration and $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ are hyperparameters depending on the system properties. The LSTM model is designed to predict the water level at the next time step based on the current water level and control input. The current water level is obtained from sensors, which are assumed to be subject to sensor noise. Formally, we define the input-output relationship of the LSTM as $x(t)=(u(t),h(t)+w(t))$ and $y(t)=h(t+1)$, where $w\sim\mathcal{N}(0,{0.01}^{2})$.
Baseline
To evaluate our proposed recovery time estimation method and adjustment technique, we prepared a baseline model: a standard model trained without considering resilience. As shown in Figure 3, the baseline model closely approximates the target system for inference.
@@ 7.2.2 Results and Analysis
First, regarding (A), although the result in Figure 4 is not contradict Theorem 5.3 in the sense of weakly decreasing, the magnitude of this decay was notably smaller than the results in Section 5.1.
Next, concerning (B), we observed a trade-off between recovery times and MAE loss when comparing the baseline model and our models, and when varying $\varepsilon$ of our model, as illustrated in Figure 4. Specifically, as $\varepsilon$ increases from 0.10.10.1 to 0.50.50.5, we observe a trend of increasing MAE loss and decreasing recovery time. This trend aligns with our theoretical predictions in Section 6.2.
@@
Figure 3: Baseline of two-tank system. Top: The blue line depicts the resulting tank water level $h_{1}$ while the orange line illustrates the LSTM prediction. Bottom: control input signal. We can confirm that the baseline model closely approximates the target system.
Figure 4: Trade-off controlled by ${\bm{\varepsilon}}$. The red dot represents the baseline value. For each $\varepsilon$ from 0.10.10.1 to 0.50.50.5 in 0.050.050.05 increments value, gray, orange, and blue dots represent MAE and $T_{R},\bar{T}_{R}(0)$, and $\bar{T}_{R}(20)$ pairs, respectively, where MAE is the loss calculated on test data. The figure shows a trade-off between recovery times and MAE when comparing the baseline and our models, and when varying $\varepsilon$ of our model.
@@ 8 Discussion
  • Theoretical and data-independent guarantees may not always reflect actual data distributions. Therefore, we strongly recommend complementing our approach with data-driven quality assurance methods for practical implementation. Also, be aware that applying our method to an unstable system can be counterproductive.
@@
  • Although our current research focused on the resilience of LSTM modules in control systems, comprehensive quality assurance for control systems remains a challenging issue. Refining studies that have demonstrated the relationship between LSTM stability and control system stability (Terzi et al.,, 2021) could be a key approach. By more precisely analyzing the calculations used to prove control system stability in these studies, it may be possible to quantitatively determine the convergence time of the state of the system to the target value.
Acknowledgment
We are grateful to Professor Tomoyuki Tanaka of Yokohama National University for helpful discussions and comments on the manuscript.
Yoshihara Sota was financially supported by JST SPRING, Grant Number JPMJSP2125. He would like to take this opportunity to thank the “THERS Make New Standards Program for the Next Generation Researchers.”
Appendix
@@ Appendix A Other Related Work
Finite time stability (FTS) (Bhat and Bernstein,, 2000) rigorously discusses convergence time, characterized by a settling-time function. Finite-time input-to-state stability (FTISS), combining FTS and ISS, enables robust control against disturbances (He et al.,, 2022). Recent studies (Liang and Liang,, 2023; Mo et al.,, 2023) propose Lyapunov functions satisfying FTISS, eliminating the need to directly consider the settling-time function.
However, deriving sufficient FTS conditions involves strict model constraints, established only for specific models like Hopfield networks (Michalak and Nowakowski,, 2017) and Clifford-valued RNNs (Shen and Meng,, 2023), but not for conventional RNNs or LSTMs. While achieving FTS for LSTMs would theoretically allow for precise recovery time evaluation, we argue that from a practical standpoint, considering numerical errors and rounding, it suffices for output differences to fall below a threshold rather than strictly reaching equilibrium. Thus, we opted for $\delta$ISS instead of FTS.
@@ Appendix B Proofs
From the definition in Section 4.1, $0\leq G^{(l)}_{*}(\eta^{(l)}(k-1)),G^{(l)}_{c}(\eta^{(l)}(k-1))$. Since $\sigma$ and $\phi$ are strictly monotonically increasing functions, we have $1/2=\sigma(0)\leq\overline{\sigma}^{(l)}_{*}(k)<1$ and $0=\phi(0)\leq\overline{\phi}^{(l)}_{c}(k)<1$. Therefore, $0\leq\overline{\sigma}^{(l)}_{i}(k)\overline{\phi}^{(l)}_{c}(k)$ and $2\leq 1/(1-\overline{\sigma}^{(l)}_{f}(k))$ are obtained, then,
$0=\phi\left(0\times 2\right)\frac{1}{2}\leq\phi\left(\frac{\overline{\sigma}^{(l)}_{i}(k)\bar{\phi}_{c}(k)}{1-\overline{\sigma}^{(l)}_{f}(k)}\right)\overline{\sigma}^{(l)}_{o}(k)=\eta(k).$ (38)
@@
Firstly, consider the case of $k=0$. From the above inequalities, and the fact that $\sigma(x)\to\infty$ as $x\to\infty$, and that $0<\overline{\sigma}^{(l)}_{o}(0)<1$, we have
@@
Since $\sigma$ and $\phi$ are strictly monotonically increasing functions, the following hold:
$\overline{\sigma}^{(l)}_{*}(1)$ $=\sigma(G_{*}(\eta^{(l)}(0)))\leq\sigma(G_{*}(\eta^{(l)}(-1)))=\overline{\sigma}^{(l)}_{k}(0),$ (45)
$\overline{\phi}^{(l)}_{c}(1)$ $=\phi(G_{c}(\eta^{(l)}(0)))\leq\phi(G_{c}(\eta^{(l)}(-1)))=\overline{\phi}^{(l)}_{c}(0).$ (46)
Therefore, $\overline{\sigma}^{(l)}_{*}(1)\leq\overline{\sigma}^{(l)}_{*}(1)$ and $\overline{\phi}^{(l)}_{c}(1)\leq\overline{\phi}^{(l)}_{c}(0)$. Applying these inequality for $\overline{c}^{(l)}(1)$ and $\eta^{(l)}(1)$, we have
@@
$\quad+\sigma(W^{(l)}_{i}x^{(l)}(t)+U^{(l)}_{i}h^{(l)}(t)+b^{(l)}_{i})_{(j)}\times\phi(W^{(l)}_{c}x^{(l)}(t)+U^{(l)}_{c}h^{(l)}(t)+b^{(l)}_{c})_{(j)}$ (58)
@@ B.2 Approximation of $\eta^{(l)}(\infty)$
Definition B.3 (ISS, Sontag, 1990; Jiang and Wang, 2001).
@@
Remarking the invariant set in Proposition 4.1, the sufficient condition for ISS of LSTM is relaxed (in the case of $k\geq 1$) than (Terzi et al.,, 2021, Theorem 1, which is the case of $k=0$) as follows.
@@
From (LABEL:c^l-LSTM) and the fact Lipschitz constant of $\phi$ is 1, we have
@@
for any $k\in\mathbb{Z}_{\geq 0}$ and $l\in[L]$. Similarly, (3c) imply that
@@
for $l=1,\ldots,L-1$, where $\llcorner A\lrcorner$ denotes the lower horizontal vector $(A_{(2,1)},A_{(2,2)})$ for a matrix $A=\begin{bmatrix}A_{(1,1)}&A_{(1,2)}\\ A_{(2,1)}&A_{(2,2)}\end{bmatrix}$. Applying the relation (3e) and inequality (100) to the second term of (99) repeatedly, we estimate that
@@
$\quad+\left\|{\sum_{\tau_{L}=1}^{t}A_{s}^{(L)}(k)^{t-\tau_{L}}a_{x}^{(L)}(k)\sum_{\tau_{L-1}=1}^{\tau_{L}}\llcorner A_{s}^{(L-1)}(k)^{\tau_{L}-\tau_{L-1}}\lrcorner a_{x}^{(L-1)}(k)\left\|{h^{(L-2)}(\tau_{L-1})}\right\|}\right\|$ (103)
@@
where $M_{L}(k),M^{\prime}_{L}(k)$, and $M^{\prime\prime}_{L}(k)$ are constants independent of $t$ and $0<\rho_{L}(k),\rho^{\prime}_{L}(k)<1$. Therefore, if $\rho(A_{s}^{(l)}(k))\in(0,1)$ for any $l\in[L]$, the multi-layer LSTM is ISS.
We note that $\mu^{(l)}$ can be expressed by direct calculation as
@@
$-(A^{(l)}_{s}(k)-A^{(l)}_{s}(k+1))v=\rho(A^{{(l)}}_{s}(k+1))v-A^{(l)}_{s}(k)v=(\rho(A^{{(l)}}_{s}(k+1))I-A_{s}^{(l)}(k))v$ (128)
is non-positive. Since all eigenvalues of $A^{{(l)}}_{s}(k)$ less than $\rho(A^{{(l)}}_{s}(k+1))$, then $\rho(A^{{(l)}}_{s}(k+1))I-A^{(l)}_{s}(k)$ is a regular matrix, and which inverse matrix can be expanded as
@@
Therefore, $(\rho(A^{{(l)}}_{s}(k+1))I-A^{(l)}_{s}(k))^{-1}$ is positive matrix. Hence, $v$ is non-positive from (128). Since $v$ is also a non-negative vector, $v=0$. This contradicts to $v$ is an eigenvector. Thus, we conclude that $\rho(A^{{(l)}}_{s}(k))\geq\rho(A^{{(l)}}_{s}(k+1))$. ∎
@@
From (LABEL:c^l-LSTM) and (3c), we have
@@
Since the Lipschitz constant of $\sigma$ is 1/41/41 / 4, we obtain that
@@
$\begin{pmatrix}\left\|{c_{1}^{(l)}(t+1)-c_{2}^{(l)}(t+1)}\right\|\\ \left\|{h_{1}^{(l)}(t+1)-h_{2}^{(l)}(t+1)}\right\|\end{pmatrix}$ $\leq\begin{bmatrix}\overline{\sigma}_{f}^{(l)}(k)&\alpha_{s}^{(l)}(k)\\ \overline{\sigma}_{o}^{(l)}(k)\overline{\sigma}_{f}^{(l)}&\alpha_{s}^{(l)}(k)\overline{\sigma}_{o}^{(l)}(k)+\frac{1}{4}\phi\left(\overline{c}^{(l)}(k)\right)\left\|{U_{o}^{(l)}}\right\|\end{bmatrix}\begin{pmatrix}\left\|{c_{1}^{(l)}(t)-c_{2}^{(l)}(t)}\right\|\\ \left\|{h_{1}^{(l)}(t)-h_{2}^{(l)}(t)}\right\|\end{pmatrix}$ (156)
@@
where $\alpha_{s}^{(l)}(k)=\frac{1}{4}\overline{c}^{(l)}(k)\|U_{f}^{(l)}\|+\overline{\sigma}_{i}^{(l)}(k)\|U_{c}^{(l)}\|+\frac{1}{4}\overline{\phi}_{c}^{(l)}(k)\|U_{i}^{(l)}\|$ and $\alpha_{x}^{(l)}(k)=\frac{1}{4}\overline{c}^{(l)}(k)\|W_{f}^{(l)}\|+\overline{\sigma}_{i}^{(l)}(k)\|W_{c}^{(l)}\|+\frac{1}{4}\overline{\phi}_{c}^{(l)}(k)\|W_{i}^{(l)}\|$.
@@
$\left\|{s_{1}(t)-s_{2}(t)}\right\|$ $\leq\mu\lambda_{1}(A_{s}(k))^{t}\left\|{s_{1}(0)-s_{2}(0)}\right\|+\left\|{(I-A_{s}(k)^{t})(I-A_{s}(k))^{-1}a_{x}(k)}\right\|\left\|{x_{1}-x_{2}}\right\|_{2,\infty}$ (159)
@@
$\left\|{s_{1}^{(L)}(t)-s_{2}^{(L)}(t)}\right\|$ $\leq\mu^{(L)}\rho(A_{s}^{(L)}(k))^{t}\left\|{s_{1}^{(L)}(0)-s_{2}^{(L)}(0)}\right\|$ (160)